# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 5.1. Analytic continuation
(5.1.6) $\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in In_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$
@@
First consider the function $f_{\beta,\delta,k}(q_{k,+})$ defined as
$\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}(-1)^{(r-1)e_{k}}}{\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})}$
$=$ $\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\left(\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}(-1)^{(n+r-1)e_{k}+\sum_{j=1}^{n}\beta\cdot M_{j}}\prod_{j=1}^{n}\sin((x_{\delta_{k}}-z_{j})/2\mathbbm{i})}{\pi^{n}}\right.$
$\cdot\left.\prod_{j=1}^{n}\frac{\Gamma((x_{\delta_{k}}-z_{j})/2\pi\mathbbm{i}+\beta\cdot M_{j}+e_{k})}{\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})}\right)$
@@
$\left[\prod_{j=1}^{n}(\partial_{k}-x_{j}/2\pi\mathbbm{i}+\beta\cdot L_{j})-q_{k,+}(-1)^{n-r+1}\prod_{j=1}^{n}(\partial_{k}-z_{j}/2\pi\mathbbm{i}+\beta\cdot M_{j})\right]\Phi(q_{k,+})=0,$
where $\partial_{k}=q_{k,+}\frac{\partial}{\partial q_{k,+}}$. This differential equation has singularities at $0,\infty,$ and $(-1)^{n-r+1}$. Following (A.0.3), our path of analytic continuation in the $\log(q_{k,+})$ plane goes through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi\mathbbm{i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0.
Consider the product
$\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})$
viewed as an analytic function of the variables $\log(q_{1,+}),\ldots,\log(q_{r,+})$. We see that it may be analytically continued everywhere except along the poles:
(5.1.7) $\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}.$
@@
$\hat{\Delta}=\prod_{k=1}^{r}\left(\prod_{i<k}(-\partial_{k}+\partial_{i})\right).$
@@
$K^{+}_{\beta,\delta}=\hat{\Delta}\left(\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})\right),$
and therefore $K^{+}_{\beta,\delta}$ has the same poles at (5.1.7).
@@
Figure 1.
Let $\gamma_{1}$ denote the path that keeps $\log(q_{2,+}),\ldots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ from Figure 1 in $\log(q_{1,+})$. Let $\gamma_{2}$ denote the path that keeps $\log(q_{1,+})$ constant (equal to $\infty$), keeps $\log(q_{3,+}),\ldots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ in $\log(q_{2,+})$. Define $\gamma_{k}$ for $k\leq r$ similarly. The analytic continuation for $K^{+}_{\beta,\delta}$ computed in (5.1.4) involved the concatenated path
$\gamma_{1}\star\cdots\star\gamma_{r}.$
We must show that this composition of paths is homotopy equivalent to a path $\hat{\gamma}$ which follows $\gamma$ in the plane $\log(q_{1,+})=\cdots=\log(q_{r,+})$ via a homotopy which is contained in
$\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{r}.$
@@
  • for all points on the path such that $\mathfrak{Re}(\log(q_{1,+}))\geq 0$ we let $\mathfrak{Im}(\log(q_{1,+}))=(n-r)\pi\mathbbm{i}$;
@@
  • for all points on the path such that $\mathfrak{Re}(\log(q_{2,+}))\leq 0$ we let $\mathfrak{Im}(\log(q_{2,+}))=(n-r)\pi\mathbbm{i}$;
@@
Figure 2. The projections of the paths $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ to the $\log(q_{1,+})$ and $\log(q_{2,+})$ planes respectively.
It is clear that $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ is homotopic to $\gamma_{1}\star\gamma_{2}$ in $\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{r}.$ By the construction, the imaginary parts $\mathfrak{Im}(\log(q_{1,+}))$ and $\mathfrak{Im}(\log(q_{2,+}))$ are equal for all points on the path $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$.
Finally, we homotope $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ further by modifying the real part of the path so that $\mathfrak{Re}(\log(q_{1,+}))=\mathfrak{Re}(\log(q_{2,+}))$ for all points on the path (see Figure 3). This second homotopy is also contained entirely in $\mathbb{C}^{2}\setminus\{\log(q_{k,+})=(n-2+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{2}.$
The resulting path $\hat{\gamma}$ from $(\log(q_{1,+}),\log(q_{2,+}))=(-\infty,-\infty)$ to (,)($\infty$,$\infty$)( ∞ , ∞ ) lies entirely in the plane $\log(q_{1,+})=\log(q_{2,+})$ as desired.
@@
Figure 3. The (real part of the) original path of analytic continuation $\gamma_{1}\star\gamma_{2}$ (in purple) and the new path $\hat{\gamma}$ (in blue).
5.2. Simplification
After specializing the Novikov parameters, we observe that $H^{{\pm}}_{\beta,\delta}|_{q_{1,\pm}=\cdots=q_{r,\pm}}=H^{{\pm}}_{\beta,\delta^{\prime}}|_{q_{1,\pm}=\cdots=q_{r,\pm}}$ whenever $\delta$ and $\delta^{\prime}$ are equal up to a permutation. This allows for a dramatic simplification to the formula (5.1.6). The simplification relies on the following identity.
@@ 5.2. Simplification
We have the following identity of (bi)-anti-symmetric functions:
(5.2.1) $\prod_{i=1}^{r}\prod_{l<i}\left((z_{i}-z_{l})(x_{l}-x_{i})\right)=\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l}).$
@@
Denote the left and right hand sides of (5.2.1) by $L(\underline{x},\underline{z})$ and $R(\underline{x},\underline{z})$ respectively.
We first claim that $R(\underline{x},\underline{z})$ is divisible by $(z_{i}-z_{l})$ for $i\neq l$. Consider the substitution $z_{i}=z_{l}=z$. Group the elements $\sigma\in S_{r}$ in pairs by grouping $\sigma$ with $\tilde{\sigma}=\sigma\circ\tau_{il}$ where $\tau_{il}$ is the transposition $(il)$. The elements $\tilde{\sigma}$ and $\sigma$ have opposite signs. Since $\sigma(l)=\tilde{\sigma}(i)$ and $\sigma(i)=\tilde{\sigma}(l)$, we see that
$\left(\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l})+\operatorname{sgn}(\tilde{\sigma})\prod_{l=1}^{r}\prod_{j\neq\tilde{\sigma}(l)}(x_{j}-z_{l})\right)|_{z_{i}=z_{l}=z}=0,$
@@
For $\delta^{-}:\{1,\ldots,r\}\to\{1,\ldots,n\}$ an injective function, let $S_{r}\cdot\delta^{-}$ denote the set of all possible functions obtained by pre-composing $\delta^{-}$ with a permutation of $\{1,\ldots,r\}.$ The identity
(5.2.2) $\sum_{f^{-}\in S_{r}\cdot\delta^{-}}C^{H}_{f^{-},\delta^{+}}=C_{\delta^{-},\delta^{+}}.$
holds, where $C_{\delta^{-},\delta^{+}}$ was defined in (2.3.4).
@@
By reindexing the sets $\{x_{1},\ldots,x_{n}\}$ and $\{z_{1},\ldots,z_{n}\}$, we may assume without loss of generality that $\delta^{-}=\delta^{+}=\{1,\ldots,r\}$. Then in this case (5.2.2) simplifies to the equation
(5.2.3) $\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq n}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})}{\sin((x_{k}-x_{i})/2\mathbbm{i})}\right)$
$=$ $\prod_{k=1}^{r}\prod_{j>r}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{k}-z_{j})/2\mathbbm{i})}$
@@
(5.2.4) $\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})}{\sin((x_{k}-x_{i})/2\mathbbm{i})}\right)=1.$
@@
$\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{1}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})\right)$
$=$ $\prod_{k=1}^{r}\prod_{j>k}\frac{1}{\sin((z_{\sigma(k)}-z_{\sigma(j)})/2\mathbbm{i})}$
$=$ $\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{j<k}\frac{1}{\sin((z_{j}-z_{k})/2\mathbbm{i})}$
and (5.2.4) is then equivalent to the equality
(5.2.5) $\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\sin((x_{k}-z_{j})/2\mathbbm{i})=\prod_{k=1}^{r}\prod_{i<k}\sin((z_{i}-z_{k})/2\mathbbm{i})\sin((x_{k}-x_{i})/2\mathbbm{i})$
@@
$\sin((a-b)/2\mathbbm{i})=\frac{e^{a/2+b/2}(e^{a}-e^{b})}{2\mathbbm{i}},$
under the change of variables $X_{i}=e^{x_{i}},Z_{i}=e^{z_{i}}$, (5.2.5) becomes
(5.2.6) $\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}(X_{k}-Z_{j})=\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\prod_{i<k}(Z_{i}-Z_{k})(X_{k}-X_{i}).$
Lemma 5.2 then gives the desired equality.
5.3. Main theorem
Combining (5.1.6) with (5.2.2), we conclude that
(5.3.1) $\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in D_{-}}C_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$
where the path of analytic continuation is along $\hat{\gamma}$ from Proposition 5.1 and the sum is now over all unordered size $r$ subsets of $\{1,\ldots,n\}$.
Let $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined in (2.3.5). Then (5.3.1) implies that:
(5.3.2) $\mathbb{U}_{H}H_{X_{-},\beta}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}=\widetilde{H_{X_{+},\beta}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}$

### v2

@@ 5.1 Analytic continuation
$\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in{\rm In}_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$ (5.6)
@@
First, consider the function $f_{\beta,\delta,k}(q_{k,+})$ defined as
$\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi{\rm i}+e_{k})}(-1)^{(r-1)e_{k}}}{\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})}$
$\qquad{}=\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\left(\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi{\rm i}+e_{k})}(-1)^{(n+r-1)e_{k}+\sum_{j=1}^{n}\beta\cdot M_{j}}\prod_{j=1}^{n}\sin((x_{\delta_{k}}-z_{j})/2{\rm i})}{\pi^{n}}\right.$
[…]
@@
$\left[\prod_{j=1}^{n}(\partial_{k}-x_{j}/2\pi{\rm i}+\beta\cdot L_{j})-q_{k,+}(-1)^{n-r+1}\prod_{j=1}^{n}(\partial_{k}-z_{j}/2\pi{\rm i}+\beta\cdot M_{j})\right]\Phi(q_{k,+})=0,$
where $\partial_{k}=q_{k,+}\frac{\partial}{\partial q_{k,+}}$. This differential equation has singularities at 00, $\infty$,∞ , and $(-1)^{n-r+1}$. Following (A.3), our path of analytic continuation in the $\log(q_{k,+})$ plane goes through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi{\rm i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0. Consider the product $\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})$ viewed as an analytic function of the variables $\log(q_{1,+}),\allowbreak\dots,\log(q_{r,+})$. We see that it may be analytically continued everywhere except along the poles:
$\log(q_{k,+})=(n-r+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}.$ (5.7)
@@
$\hat{\Delta}=\prod_{k=1}^{r}\biggl{(}\prod_{i<k}(-\partial_{k}+\partial_{i})\biggr{)}.$
@@
$K^{+}_{\beta,\delta}=\hat{\Delta}\Biggl{(}\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})\Biggr{)},$
and therefore $K^{+}_{\beta,\delta}$ has the same poles at (5.7).
@@
Figure 1:
Let $\gamma_{1}$ denote the path that keeps $\log(q_{2,+}),\dots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ from Figure 1 in $\log(q_{1,+})$. Let $\gamma_{2}$ denote the path that keeps $\log(q_{1,+})$ constant (equal to $\infty$), keeps $\log(q_{3,+}),\dots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ in $\log(q_{2,+})$. Define $\gamma_{k}$ for $k\leq r$ similarly. The analytic continuation for $K^{+}_{\beta,\delta}$ computed in (5.4) involved the concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$. We must show that this composition of paths is homotopy equivalent to a path $\hat{\gamma}$ which follows $\gamma$ in the plane $\log(q_{1,+})=\cdots=\log(q_{r,+})$ via a homotopy which is contained in
$\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}\}_{k=1}^{r}.$
@@
  • for all points on the path such that $\mathfrak{Re}(\log(q_{1,+}))\geq 0$ we let $\mathfrak{Im}(\log(q_{1,+}))=(n-r)\pi{\rm i}$;
@@
  • for all points on the path such that $\mathfrak{Re}(\log(q_{2,+}))\leq 0$ we let $\mathfrak{Im}(\log(q_{2,+}))=(n-r)\pi{\rm i}$;
@@
Figure 2: The projections of the paths $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ to the $\log(q_{1,+})$ and $\log(q_{2,+})$ planes, respectively.
It is clear that $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ is homotopic to $\gamma_{1}\star\gamma_{2}$ in $\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}\}_{k=1}^{r}.$ By the construction, the imaginary parts $\mathfrak{Im}(\log(q_{1,+}))$ and $\mathfrak{Im}(\log(q_{2,+}))$ are equal for all points on the path $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$.
Finally, we homotope $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ further by modifying the real part of the path so that we have $\mathfrak{Re}(\log(q_{1,+}))=\mathfrak{Re}(\log(q_{2,+}))$ for all points on the path (see Figure 3). This second homotopy is also contained entirely in $\mathbb{C}^{2}\setminus\{\log(q_{k,+})=(n-2+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}\}_{k=1}^{2}.$
The resulting path $\hat{\gamma}$ from $(\log(q_{1,+}),\log(q_{2,+}))=(-\infty,-\infty)$ to (,)($\infty$,$\infty$)( ∞ , ∞ ) lies entirely in the plane $\log(q_{1,+})=\log(q_{2,+})$ as desired. ∎
@@
Figure 3: The (real part of the) original path of analytic continuation $\gamma_{1}\star\gamma_{2}$ (in purple) and the new path $\hat{\gamma}$ (in blue).
5.2 Simplification
After specializing the Novikov parameters, we observe that $H^{{\pm}}_{\beta,\delta}|_{q_{1,\pm}=\cdots=q_{r,\pm}}=H^{{\pm}}_{\beta,\delta^{\prime}}|_{q_{1,\pm}=\cdots=q_{r,\pm}}$ whenever $\delta$ and $\delta^{\prime}$ are equal up to a permutation. This allows for a dramatic simplification to the formula (5.6). The simplification relies on the following identity.
@@ 5.2 Simplification
We have the following identity of ((((bi))))-anti-symmetric functions:
$\prod_{i=1}^{r}\prod_{l<i}((z_{i}-z_{l})(x_{l}-x_{i}))=\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l}).$ (5.8)
@@
Denote the left- and right-hand sides of (5.8) by $L(\underline{x},\underline{z})$ and $R(\underline{x},\underline{z})$, respectively. We first claim that $R(\underline{x},\underline{z})$ is divisible by $(z_{i}-z_{l})$ for $i\neq l$. Consider the substitution $z_{i}=z_{l}=z$. Group the elements $\sigma\in S_{r}$ in pairs by grouping $\sigma$ with $\tilde{\sigma}=\sigma\circ\tau_{il}$ where $\tau_{il}$ is the transposition $(il)$. The elements $\tilde{\sigma}$ and $\sigma$ have opposite signs. Since $\sigma(l)=\tilde{\sigma}(i)$ and $\sigma(i)=\tilde{\sigma}(l)$, we see that
$\Biggl{(}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l})+\operatorname{sgn}(\tilde{\sigma})\prod_{l=1}^{r}\prod_{j\neq\tilde{\sigma}(l)}(x_{j}-z_{l})\Biggr{)}\bigg{|}_{z_{i}=z_{l}=z}=0,$
@@
For $\delta^{-}\colon\{1,\dots,r\}\to\{1,\dots,n\}$ an injective function, let $S_{r}\cdot\delta^{-}$ denote the set of all possible functions obtained by pre-composing $\delta^{-}$ with a permutation of $\{1,\dots,r\}.$ The identity
$\sum_{f^{-}\in S_{r}\cdot\delta^{-}}C^{H}_{f^{-},\delta^{+}}=C_{\delta^{-},\delta^{+}}$ (5.9)
holds, where $C_{\delta^{-},\delta^{+}}$ was defined in (2.7).
@@
By reindexing the sets $\{x_{1},\dots,x_{n}\}$ and $\{z_{1},\dots,z_{n}\}$, we may assume without loss of generality that $\delta^{-}=\delta^{+}=\{1,\dots,r\}$. Then in this case, (5.9) simplifies to the equation
$\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\Biggl{(}\prod_{\stackrel{{\scriptstyle 1\leq j\leq n}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2{\rm i})}{\sin((z_{\sigma(k)}-z_{j})/2{\rm i})}\prod_{i<k}\frac{\sin\bigl{(}\bigl{(}z_{\sigma(k)}-z_{\sigma(i)}\bigr{)}/2{\rm i}\bigr{)}}{\sin((x_{k}-x_{i})/2{\rm i})}\Biggr{)}$
$\qquad{}=\prod_{k=1}^{r}\prod_{j>r}\frac{\sin((x_{k}-z_{j})/2{\rm i})}{\sin((z_{k}-z_{j})/2{\rm i})}$
@@
$\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\Biggl{(}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2{\rm i})}{\sin((z_{\sigma(k)}-z_{j})/2{\rm i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2{\rm i})}{\sin((x_{k}-x_{i})/2{\rm i})}\Biggr{)}=1.$ (5.10)
@@
$\prod_{k=1}^{r}\Biggl{(}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{1}{\sin((z_{\sigma(k)}-z_{j})/2{\rm i})}\prod_{i<k}\sin((z_{\sigma(k)}-z_{\sigma(i)})/2{\rm i})\Biggr{)}$
$\qquad{}=\prod_{k=1}^{r}\prod_{j>k}\frac{1}{\sin((z_{\sigma(k)}-z_{\sigma(j)})/2{\rm i})}=\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{j<k}\frac{1}{\sin((z_{j}-z_{k})/2{\rm i})}$
and (5.10) is then equivalent to the equality
$\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\sin((x_{k}-z_{j})/2{\rm i})=\prod_{k=1}^{r}\prod_{i<k}\sin((z_{i}-z_{k})/2{\rm i})\sin((x_{k}-x_{i})/2{\rm i}).$ (5.11)
@@
$\sin((a-b)/2{\rm i})=\frac{{\rm e}^{a/2+b/2}\bigl{(}{\rm e}^{a}-{\rm e}^{b}\bigr{)}}{2{\rm i}},$
under the change of variables $X_{i}={\rm e}^{x_{i}}$, $Z_{i}={\rm e}^{z_{i}}$, (5.11) becomes
$\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}(X_{k}-Z_{j})=\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\prod_{i<k}(Z_{i}-Z_{k})(X_{k}-X_{i}).$
Lemma 5.2 then gives the desired equality. ∎
5.3 Main theorem
Combining (5.6) with (5.9), we conclude that
$\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in D_{-}}C_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}},$ (5.12)
where the path of analytic continuation is along $\hat{\gamma}$ from Proposition 5.1 and the sum is now over all unordered size $r$ subsets of $\{1,\dots,n\}$.
Let $\mathbb{U}_{H}\colon H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined in (2.8). Then (5.12) implies that
$\mathbb{U}_{H}H_{X_{-},\beta}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}=\widetilde{H_{X_{+},\beta}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}$ (5.13)